Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability Problem

Expert replies
by kidboc1 » Sat Sep 20, 2008 4:40 pm
Out of a group of four men and six women, three people will be randomly chosen to participate in a marketing survey. What is the probability that at least two of the people chosen is a man?

thanks
Join the discussion
Source: — Problem Solving |

by varun nagpal » Sat Sep 20, 2008 8:28 pm
ans is 5/12

probability to choose 0 man ot 1 man

which is 3/6 (all woman) + 1/4*2/6 (one man and 2 woman) = 7/12

1 - 7/12 = 5/12 prob of choosing atleast 2 man
Join the discussion

by varun nagpal » Sat Sep 20, 2008 8:29 pm
ans is 5/12

probability to choose 0 man ot 1 man

which is 3/6 (all woman) + 1/4*2/6 (one man and 2 woman) = 7/12

1 - 7/12 = 5/12 prob of choosing atleast 2 man
Join the discussion

by ssy » Mon Sep 22, 2008 1:20 am
Kidboc1 - What is the answer? Please provide answer options.

Using a similar method to Varun, I got ans=1/3.

The probabily of it being all woman is: [(6/10)(5/9)(4/8)] = 1/6

The probabilty of only having 1 man out of 3 (this is probably the trickiest part) is:

If man is chosen first:[(4/10)+(6/9)+(5 / 8)] = 1/6
If man is chosen second: [(6/10)+(4/9)+(5 / 8)] = 1/6
If man is chosen third: [(6/10)+(5/9)+(4 / 8)] = 1/6

Probability of NOT having at least 2 man = 1/6+[(1/6)+(1/6)+(1/6)] = (1/6)+(1/2) = 4/6 = 2/3

Probability of having at least 2 men = 1-(2/3) = 1/3
Last edited by ssy on Mon Sep 22, 2008 1:24 am, edited 1 time in total.
Join the discussion

by ssy » Mon Sep 22, 2008 1:23 am
I am not quite sure why my number 8 have turned into smiley faces in my post above!

Please replace all smiley faces with the number 8.
Join the discussion

by vr4indian » Mon Sep 22, 2008 2:41 am
I think answer is 1/3

Part 1:
Total way to select 3 people from 10 = 10c3 = 120

Part 2:
We have 2 men and 1 woman
4c2.6c1 = 36 possibilities.

We can have 3 men team only
4c3.6c0 = 4 possibilities.

So total way = 36 +4 = 40

Prob . 40 / 120 = 1/3

Hope i am clear and its very easy method .

Thanks
Vishal Shah
Join the discussion

re:

by kidboc1 » Mon Sep 22, 2008 1:48 pm
Thank you all for your responses. 1/3 is the correct answer.
Join the discussion

re:

by kidboc1 » Mon Sep 22, 2008 2:45 pm
Thank you all for your responses. 1/3 is the correct answer.
Join the discussion

by varun nagpal » Tue Sep 23, 2008 9:36 am
agree 1/3 is correct ans my method is flaw
Join the discussion

by vanesslondon » Tue Sep 23, 2008 1:35 pm
Hi guys, thanks for all your input, I did this a different way. I calculated all the possibilities for men being chosen.


MMM

4/10 * 3/9 * 2/8

MWM

4/10 * 6/9 * 3/8

MMW

4/10 * 3/9 * 6/8

WMM

6/10 * 4/9 * 3/8

I then added the answers and got 1/3
Join the discussion

by deepamohn » Wed Oct 08, 2008 11:27 am
Using Combination.

Result is ( 4c3+4c2 . 6c1 ) / 10c3 => 1/3 :D
Join the discussion